Determine the indicated probability for a binomial experiment with the given number of trials n and the given success probability p. n= 13,p= 0.7, /{Fewer than 4) A) 0.0007 B) 0.0001 C) 0.9993 D) 0.0040
Answer #1

solution

Given that ,

p = 0.7

1 - p = 1 - 0.7= 0.3

n = 13

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X < 4) = P(X=0) +P(X=1)+P(X=2)+P(X=3)

= (13 C 0) * 0.70 * (0.3)13+(13 C 1) * 0.71 * (0.3)12+(13 C 2) * 0.72 * (0.3)11+(13 C 3) * 0.73 * (0.3)10

probability =0.0007